EVA Report – February 25th

Crew 275 EVA Report 25Feb2023

EVA # 12

Author: Alice Chapiron, EVA Leader and Crew Scientist

Purpose of EVA: Testing the MetMet and explori0.ng the area

Start time: 9:03

End time: 11:51


09:03: Airlock closed.

09:10: Rovers Curiosity and Spirit unplugged. We then left from the Hab at 09:14.

09:16: Rovers parked near the atmospheric instruments.

09:58: Data retrieved and batteries changed for the Purple Air, LOAC, Mega-Ares, and Weather Station. The EVA crew is about to go back to the rovers to go to their target location (Kissing Camel Ridge W).

10:15: Rovers are parked at Kissing Camel Ridge, the EVA Leader pass on channel 2 in order to hear her EVA crewmates, then the EVA Crew begins to walk towards their destination for sample collection.

10:34: The EVA Crew starts drawing the area where they will take the rock samples.

10:43: The EVA Crew begins to collect samples and to do the related measurements with the MetMet.

11:04: The EVA Leader had some problems hearing the HabCom with her radio, the Crew Commander took the communications.

11:25: The EVA Crew is tidying up their material.

11:30: EVA Crew back to the Rover.

11:42: Arrival at the Hab.

11:51: End of the EVA.

Destination: Kissing Camel Ridge W

Coordinates: N4249500 E518000

Participants: EVA Leader: Alice Chapiron (Crew Scientist), Jérémy Rabineau (Crew Commander), Adrien Tison (Crew Botanist), Marie Delaroche (Crew Journalist)

HabCom: Alexandre Vinas (Crew Astronomer)

Road(s) and routes per MDRS Map: Cow Dung Road until Robert’s Rock Garden then we will walk to a maximum of 400m from the rovers

Mode of travel: Driving using Curiosity and Spirit and walking

GreenHab Report – February 25th

GreenHab Officer: Adrien Tison

Environmental control: Heater

Average Temperature: 83.7 °F

Average humidity: 28 %

Hours of supplemental light: 0 hour

Daily water usage for crops: 13.61 gallons

Daily water usage for research and/or other purposes: 0.13 gallons

Water in the blue tank: 186.26 gallons

Time(s) of watering for crops: 9:30 & 13:34

Changes to crops: Planted a bean seed in a pot containing a single bean plant.

Narrative: Like yesterday, I was on an EVA this morning, so it was once again the HSO that watered and maintained plants and crops.

I went to check on it during the afternoon and decided to water them again since it was hot in there.

I also had the pleasure to see that there were even more tomatoes to harvest!

As I said in a previous report, I was waiting for a good moment to harvest them, and this moment arrived.

I saw there was a bean that was too mature to be eaten, I then decided to take the beans inside and plant two of them in two pots containing only one bean plant.

Harvest: 1.541 kg of tomatoes

Support/supplies needed: N/A

Journalist Report – February 25th

Sol 13 – Field Day for Crew 275

“And in certain houses you heard the hard clatter of a typewriter, the novelist at work; or the scratch

of a pen, the poet at work; or no sound at all, the former beachcomber at work.”

– Chapter 13 of The Martian Chronicles by Ray Bradbury

… or the clatter of a keyboard, the journalist at work! I thought that today, I might take up the beginning of this report to mention what constitutes a significant amount of my time at MDRS. I love that I get to follow and observe each crew member in their assigned work and experiments, as well as participate in their research and lead my own. I am lucky enough to be able to balance out scientific work with more “artistic” activities: taking and editing photos of my crewmates, and doing what I love most, writing.

Today there were many opportunities to take great pictures, as I participated in the first geology EVA of the mission! Alice, our EVA leader, took Adrien, Jérémy, and I to Kissing Camel Ridge, just South of the MDRS. After exploring the area for a short while, Alice gave us instructions to draw the area we explored so that we could make note of where we had collected samples. She showed us how to choose rock samples and test them with the MetMet before collecting them. This instrument allows geologists to test the magnetic susceptibility and conductivity of rocks directly in the field; this way, one can determine the value of potential samples before even beginning to collect them.

It felt great to go a bit further away from the Hab than usual and see different landscapes, and also have new EVA objectives! The latter having been achieved, we also had a lot of fun joking with each other over the radio as we were collecting the samples. Weary but happy about our outing, we brushed off our dusty knees and returned to the Hab to a great meal prepared by Corentin. Quentin tried making bread, not quite up to his expectations…

After lunch, Jérémy started taking some time to speak individually with each of us, to check on how everyone is doing now that we have almost reached mid-rotation. Quentin also started implementing the location tracking system in the Hab to test it out before it is deployed in the entire station. In the end, there will be 10 integrated circuit boards we call “anchors” dispatched everywhere in the station, and each crew member will also wear one at all times (these we call “tags”). Every 10 seconds, the anchors and tags will “communicate” with each other and the distance between them will be saved to a database. This way, we will be able to know how much time crew members spend in different types of environments (bright or dark, noisy or calm, crowded or not, etc.) thanks to the environmental sensors. Of course, this data will be anonymized!

Coupled with the physiological and cognitive data we have been collecting, the researcher with whom we are working at the Swedish Royal Institute of Technology, will be able to analyze how the architecture and environment of the station has affected our work, our health and our mental state.

We are all looking forward to an evening full of laughs, and most importantly, a restful night.

Operations Report – February 25th

Crew 275 Operations Report 25Feb2023
SOL: 13

Name of person filing report: Quentin Royer
Non-nominal systems:
• Toilet
• Spacesuit n°9
• GreenHab Carbon Monoxide alarm

Notes on non-nominal systems: None

Spirit rover used: Yes
• Hours: 209.6
• Beginning charge: 100%
• Ending charge: 68%
• Currently charging: Yes

Opportunity rover used: No
• Hours: 113.4
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Curiosity rover used: Yes
• Hours: 221.4
• Beginning charge: 100%
• Ending charge: 68%
• Currently charging: Yes

Perseverance rover used: No
• Hours: 255.8
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

General notes on rovers: Rovers were used during today’s EVA. Nothing to report.

Summary of Hab operations: The Crew conducted routine operations.
• WATER USE: 38.62 – It is the beginning of the tank, so the formula could be inaccurate
• Water: 502.11 gallons
• Static tank pipe heater: On
• Static tank heater: On
• Toilet tank emptied: Yes

Summary of internet: Hughes Net was used during the day before the opening of COMMS for email purposes, and Starlink (MDRSCampus-Guest) is used during COMMS. MDRScampus is used for experimental purpose.

Summary of suits and radios: Suits and radios were used during today’s EVA. Nothing to report.

Summary of GreenHab operations: The Crew botanist conducted routine operations, and harvested some tomatoes.
• WATER USE: 13.74 gallons
• Heater: On
• Supplemental light: No
• Harvest: 1541 g of tomatoes

Summary of ScienceDome operations: The Crew Scientist conducted routine operations, and tests for the Human Factors experiment were performed.
• Dual split: Switched off during the night

Summary of RAM operations: No operations were conducted in the RAM today.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns, and requests to Mission Support:
• The non-nominal GreenHab Carbon Monoxide detector was placed in the RAM.

Sol Summary – February 25th

Crew 275 Sol Summary Report 25Feb2023

Sol: 13

Summary Title: Field Day for Crew 275

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

After almost two weeks at the MDRS, all the crewmembers are getting more and more efficient in all their tasks. This gives a little bit more time to relax between activities, which is appreciated by everyone. Today was the first geology EVA and it was a success. It started by a first stop at the atmospheric instruments site to change the batteries and retrieve the data, before heading to Kissing Camel Ridge W. The EVA Crew had the opportunity to test an instrument to easily measure both the magnetic susceptibility and the conductivity in real field conditions.

On the campus, everyone is now very happy to be able to follow the past and current temperature, humidity, and luminosity in all the buildings. Human factors experiments have continued, while our crew journalist has started working on our mid-rotation video. We are all looking forward to seeing the results!

As we are reaching the end of this second week, I also decided to spend some time to see each crewmember individually and talk about how they felt so far. I know that the third quarter of any long-duration challenge (like a Marathon, for instance) is the most difficult. It is also the case for analog missions like ours, and I want to make sure that everyone is in the best condition for this period.

Look Ahead Plan:

Today we have changed the batteries of the atmospheric instruments. This means that tomorrow can be a real rest day without any EVAs. Next week, we would like to plan 3 additional EVAs dedicated to photogrammetry: One for the mapping of the area, one for the exploration with a 2D map, and the third one for the exploration with a 3D render. This time the target area would be Candor Chasma.

Anomalies in work: None.

Weather: Sunny, low -4°C/25°F, high 6°C/43°F.

Crew Physical Status: Optimal.

EVA: EVA #12 was successful. All the objectives have been accomplished and the collected samples will soon be analyzed.

Reports to be filed: Sol Summary, EVA Report, Operations Report, Journalist Report, GreenHab Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site